OpenStudy (anonymous):

14. Solve log 3x – log 13 = 2. Round to the nearest thousandth if necessary.

OpenStudy (mathstudent55):

Remember this rule of logs: \( \log a - \log b = \log \dfrac{a}{b} \)

OpenStudy (mathstudent55):

Then use the definition of log: \(\log_a x = y\) means \(a^y = x\).

OpenStudy (anonymous):

so 3x-13=3*/13

OpenStudy (mathstudent55):